OpenStudy (anonymous):

find the distance of the point (1,-2,3) from the plane x-y+z=5 measured parallel to the line x/2=y/3=z/-6

ganeshie8 (ganeshie8):

start by writing the equation of line parallel to the given line, and passing thru the given point

ganeshie8 (ganeshie8):

parametric form of line passing thru (1, -2, 3) and parallel to the given line : \(\large \mathbb{x = 2t+1}\) \(\large \mathbb{y = 3t-2}\) \(\large \mathbb{z = -6t+3}\)

ganeshie8 (ganeshie8):

next, find at what point this line intersects the given plane

ganeshie8 (ganeshie8):

lastly find the distance using the distance formula

OpenStudy (lastdaywork):

I'll like to modify ganeshie's solution a li'll bit. As we are working with parametric form; we only need to use the condition that the point (2t+1, 3t-2, -6t+3) satisfies the equation of plane x-y+z=5 The absolute value of 't' will be the required distance itself. PS: I am lazy ;)